# Automatic Outlier Rectification via Optimal Transport

### v1

@@ Abstract
In this paper, we propose a novel conceptual framework to detect outliers using optimal transport with a concave cost function. Conventional outlier detection approaches typically use a two-stage procedure: first, outliers are detected and removed, and then estimation is performed on the cleaned data. However, this approach does not inform outlier removal with the estimation task, leaving room for improvement. To address this limitation, we propose an automatic outlier rectification mechanism that integrates rectification and estimation within a joint optimization framework. We take the first step to utilize an optimal transport distance with a concave cost function to construct a rectification set in the space of probability distributions. Then, we select the best distribution within the rectification set to perform the estimation task. Notably, the concave cost function we introduced in this paper is the key to making our estimator effectively identify the outlier during the optimization process. We discuss the fundamental differences between our estimator and optimal transport-based distributionally robust optimization estimator. finally, we demonstrate the effectiveness and superiority of our approach over conventional approaches in extensive simulation and empirical analyses for mean estimation, least absolute regression, and the fitting of option implied volatility surfaces.
@@ 1 Introduction
The challenge of outlier removal has been documented for many years (literally centuries); see e.g. (Gergonne, 1821) and (Peirce, 1852). Yet, the outlier removal problem continues to interest practitioners and researchers alike due to the danger of distorted model estimation mentioned above. A natural family of approaches followed in the literature takes the form of “two-stage” methods, which involve an outlier removal step followed by an estimation step. Methods within this family range from rules of thumb, such as removing outliers beyond a particular threshold based on a robust measure of scale like the interquartile range (Tukey, 1977), to various model-based or parametric distributional assumption-based tests (Thompson, 1985). While the two-stage approach can be useful by separating the estimation task from the outlier removal objective, it is not without potential pitfalls. For example, an outlier detection step that relies on a specific fitted model may tend to overfit to a particular type of outliers, which can mask the effect of other outliers (Rousseeuw & Leroy, 1987). On the other hand, the outlier detection step may be overly conservative, not being informed by a particular downstream task at hand. This can lead to a significant reduction in the number of observations available for model fitting in the next step, resulting in a loss of statistical efficiency (He & Portnoy, 1992). Despite these challenges, two-stage procedures remain popular among practitioners
@@
Robust statistics, pioneered by Box (1953); Tukey (1960); Huber (1964) and others such as Hampel (1968, 1971) offers alternative approaches for obtaining statistical estimators in the presence of outliers without removing them, particularly in parametric estimation problems such as linear regression. Beyond these, one closely related approach is the minimum distance functionals-based estimator, introduced by Parr & Schucany (1980); Millar (1981); Donoho & Liu (1988a, b); Park et al. (1995); Zhu et al. (2022); Jaenada et al. (2022). This method involves projecting the corrupted distribution onto a family of distributions using a distribution discrepancy/metric and selecting the optimal estimator for the resulting distribution. It is important to note that this projection mechanism is not informed by the estimation task (such as fitting a linear regression or neural network). Thus, without additional information about the contamination model or the estimation task, it can be challenging to choose an appropriate family of distributions for the projection step, which may lead to similar limitations as those outlined earlier in practice.
@@
In this paper, we propose a novel approach to integrate both outlier detection and estimation in a joint optimization framework. A key observation is that statisticians aim to “clean” data before decisions are made; thus, ideal robust statistical estimators tend to be optimistic. To address this, we introduce the rectification set—a ball centered around the empirical (contaminated) distribution and defined by the optimal transport distance (see (Villani, 2009; Peyré et al., 2019)) in probability space. This rectification set aims to exclude potential outliers and capture the true underlying distribution, allowing us to minimize the expectation over the “best-case” scenarios, leading to a min-min-type formulation. The study in Jiang & Xie (2024) also explores the connections with min-min type problems, which concentrates on the construction of some artificially constructed rectification sets to cover certain existing statistical estimators from the robust statistics literature. However, our primary focus is on introducing a novel rectification set, which is based on the optimal transport approach with a concave cost function.
@@
To automatically detect outliers during the estimation process, one of our main contributions is the use of a concave cost function for the optimal transport distance. This function encourages “long hauls” transportation, in which the optimal transport plan moves only a portion of the data to a distant location, while leaving other parts unchanged. This strategic approach effectively repositions identified outliers closer to positions that better align with the clean data. Our novel formulation then involves minimizing the expected loss under the optimally rectified empirical distribution. This rectification is executed through the application of an optimal transport distance with a concave cost function, thereby correcting outliers to enhance performance within a fixed estimation task.
More importantly, our method distinguishes itself from distributionally robust optimization (DRO) (Ben-Tal et al., 2013; Bayraksan & Love, 2015; Wiesemann et al., 2014; Delage & Ye, 2010; Gao & Kleywegt, 2022; Blanchet & Kang, 2017; Shafieezadeh Abadeh et al., 2015; Shafieezadeh-Abadeh et al., 2019; Sinha et al., 2018; Kuhn et al., 2019) due to the distinctive correction mechanisms initiated by the robust formulation employing a min-min strategy.
@@
As we discuss in Section 2 below, the timing of error generating differs crucially between DRO and robust statistics. DRO approach employs a min-max game strategy to control the worst-case loss over potential post-decision distributional shifts. In contrast, the robust estimator acts after the pre-decision distributional contamination materializes. Thus the approach of robust statistics can be motivated as being closer to a max-min game against nature. As a consequence, in robust statistics, the adversary moves first, and therefore the statistician can be more optimistic that they can rectify the contamination applied by the nature thus motivating the min-min strategy suggested above.
@@
  1. (i)
    (Novel statistically robust estimator) We propose a new statistically robust estimator that incorporates a novel rectification set constructed using the optimal transport distance. By employing a concave cost function within the optimal transport distance, our estimator enables automatic outlier rectification. We prove that the optimal rectified distribution can be found via a finite convex program and provide a cross-validation procedure for determining the optimal uncertainty region for determining the rectification set.
  2. (ii)
    (Connection to adaptive quantile regression) For mean estimation and least absolute regression, we demonstrate that the statistically robust estimator is equivalent to an adaptive quantile (regression) estimator, with the quantile controlled by the budget parameter $\delta$. Furthermore, we prove that the optimal rectified distribution exhibits a long-haul structure that facilitates outlier detection.
  3. (iii)
    (Effectiveness in smoothing the implied volatility surface of real options) We evaluate the performance of our estimator on various tasks, including mean estimation, least absolute regression, and a real-world application involving options implied volatility surface smoothing. Our experimental results demonstrate that our estimator produces surfaces that are 30.4% smoother compared to baseline estimators, indicating success in outlier removal because the curves should be smooth based on structural financial properties. Additionally, it achieves an average reduction of 6.3% in mean average percent error (MAPE) across all estimated surfaces, providing empirical evidence of the effectiveness of our rectifying optimal transporter.
@@ 2 DRO and Robust Statistics as Adversarial Games
In this section, we summarize conceptually how robust statistics are different with DRO, see Blanchet et al. (2024) for details. To lay a solid mathematical foundation, we begin by investigating a generic stochastic optimization problem. Here, we assume that $Z$ is a random vector in space $\mathcal{Z}\subseteq\mathbb{R}^{d}$ that follows the distribution $\mathbb{P}_{\star}$. The set of feasible model parameters is denoted $\Theta$ (assumed to be finite-dimensional to simplify). Given a realization $z$ and a model parameter $\theta\in\Theta$ the corresponding loss is $\ell(\theta,z)$. A standard expected loss minimization decision rule is obtained by solving
$\min_{\theta\in\Theta}~{}\mathbb{E}_{\mathbb{P}_{\star}}[\ell(\theta,\xi)]=\int_{\Xi}\ell(\theta,\xi)~{}\mathrm{d}\mathbb{P}_{\star}(\xi).$ (2.1)
Since $\mathbb{P}_{\star}$ is generally unknown, to approximate the objective function in (2.1), we often gather $n$ independent and identically distributed (i.i.d.) samples $\{z_{i}\}_{i=1}^{n}$ from the unknown data-generating distribution $\mathbb{P}_{\star}$ and consider the empirical risk minimization counterpart,
$\min_{\theta\in\Theta}~{}\mathbb{E}_{\hat{\mathbb{P}}_{n}}[\ell(\theta,Z)]=\frac{1}{n}\sum_{i=1}^{n}\ell(\theta,z_{i}),$ (2.2)
where ^$\hat{\mathbb{P}}_{n}$ denotes the empirical measure $\frac{1}{n}\sum_{i=1}^{n}\delta_{z_{i}}$ and $\delta_{z}$ is the Dirac measure centered at $z$.
Then, we proceed to naturally introduce the general data-driven decision-making cycle as below:
pt Contaminated data-generating distribution pt Data-driven model pt Out-of-sample environment $\mathbb{P}_{\star}$$\ {\mathbb{P}^{\prime}}$$\ \ {\mathbb{P}}_{n}^{\prime}$$\ \ \widetilde{\mathbb{P}}$
@@
In the general data-driven decision-making cycle depicted in Figure 1, we usually collect $n$ i.i.d samples from the unknown data-generating distribution $\mathbb{P}^{\prime}$, which may be identical to or distinct from the clean distribution $\mathbb{P}_{\star}$. Subsequently, we make a decision (e.g., parameter estimation) based on a model, ${\mathbb{P}}_{n}^{\prime}$, built from these samples. Such a model could be parametric or non-parametric. These decisions are then put into action within the out-of-sample environment $\tilde{\mathbb{P}}$, which may or may not conform to the distribution $\mathbb{P}_{\star}$. In this general cycle, the sample average method (2.1) may lead to poor out-of-sample guarantees.
Then, we aim to distinguish between DRO and robust statistics by treating them as adversarial games. The crucial distinction lies in the timing of contamination or attacks.
(i) (DRO: $\widetilde{\mathbb{P}}\neq{\mathbb{P}}_{\star}={\mathbb{P}^{\prime}}$) The attack occurs in the post-decision stage. In this scenario, the out-of-sample environment $\tilde{\mathbb{P}}$ diverges from the data-generating distribution $\mathbb{P}_{\star}$, and no contamination occurs before the decision, implying that $\mathbb{P}_{\star}=\mathbb{P}^{\prime}$. For instance, in adversarial deployment scenarios, malicious actors can deliberately manipulate the data distribution to compromise the performance of trained models. With full access to our trained machine learning model, the adversary endeavors to create adversarial examples specifically designed to provoke errors in the model’s predictions.
To ensure good performance in terms of the optimal expected population loss over the out-of-sample distribution, the DRO framework introduces an uncertainty set $\mathcal{B}({\mathbb{P}}^{\prime}_{n})$ to encompass discrepancies between the in-of-sample-distribution ${\mathbb{P}}_{n}^{\prime}$ and the out-of-sample distribution $\tilde{\mathbb{P}}$. Subsequently, the DRO formulation minimizes the worst-case loss within this uncertainty set, thereby aiming to:
$\min_{\theta\in\Theta}\sup_{\mathbb{Q}\in\mathcal{B}(\mathbb{P}_{n}^{\prime})}\mathbb{E}_{\mathbb{Q}}[\ell(\theta,Z)].$ (2.3)
@@
$\min_{\theta\in\Theta}\min_{\mathbb{Q}\in\mathcal{R}({\mathbb{P}}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}\left[\ell(\theta,Z)\right].$ (2.4)
Again, our goal in this paper is in connection to (2.4) and thus completely different from DRO.
@@ 3 Automatic Outlier Rectification Mechanism
In this section, we introduce our primary contribution by delineating (2.4). Our estimator is crafted to incorporate outlier rectification and parameter estimation within a unified optimization framework, facilitating automatic outlier rectification.
A natural question arises from (2.4): how do we construct the rectification set in the space of probability distributions to correct the observed data set? In this paper, we employ an optimal transport distance to create a ball centered at the contaminated empirical distribution $\mathbb{P}_{n}^{\prime}$.
Definition 1 (Rectification Set).
@@
$\mathcal{R}(\mathbb{P}_{n}^{\prime})=\{\mathbb{Q}\in\mathcal{P}(\mathcal{Z}):\mathds{D}_{c}(\mathbb{Q},\mathbb{P}^{\prime}_{n})\leq\delta\},$ (3.1)
where $\delta>0$ represents a radius, and $\mathds{D}_{c}$ is a specific optimal transport distance defined in Definition 2.
Definition 2.
(Optimal Transport Distance (Peyré et al., 2019; Villani, 2009)) Suppose that $c(\cdot,\cdot):\mathcal{Z}\times\mathcal{Z}\rightarrow[0,\infty]$ is a lower semi-continuous cost function such that $c(z,z^{\prime})=0$ for all $z,z^{\prime}\in\mathcal{Z}$ satisfying $z=z^{\prime}$. The optimal transport distance between two probability measures $\mathbb{P},\mathbb{Q}\in\mathcal{P}(\mathcal{Z})$ is defined as
@@
If we consider $c(z,z^{\prime})=\|z-z^{\prime}\|^{r}$ as a metric defined on $\mathbb{R}^{d}$, where $r\in[1,\infty)$, then the distance metric $\mathds{D}_{c}^{1/r}(\mathbb{P},\mathbb{Q})$ corresponds to the $r$-th order Wasserstein distance (Villani, 2009). In this paper, we pioneer the utilization of concave cost functions, exemplified by $c(z,z^{\prime})=\|z-z^{\prime}\|^{r}$ where $r\in(0,1)$ in statistical robust estimation. We note that $\mathds{D}_{c}(\mathbb{P},\mathbb{Q})$ (as opposed to $\mathds{D}_{c}^{1/r}(\mathbb{P},\mathbb{Q})$ ) is also a metric for $r\in[0,1)$. The rationale behind selecting a concave cost function is intuitive: it promotes long hauls transportation plans, enabling outliers to be automatically moved significant distances back towards the central tendency of data distribution. This, in turn, facilitates automatic outlier rectification. Concave costs promote long hauls due to their characteristic of exhibiting decreasing marginal increases in transportation cost. In other words, if an adversary decides to transport a data point by $\|\Delta\|$ units, it becomes cheaper to continue transporting the same point by an additional $\epsilon$ distance compared to moving another point from its initial location
Remark 1.
(i) When we disregard the outer minimization concerning estimation parameters, the inner minimization problem over probability measures is closely linked to the minimum distance functionals-based estimator, initially proposed by (Donoho & Liu, 1988a). The estimator, is obtained by first projecting the corrupted distribution $\mathbb{P}^{\prime}_{n}$ onto a family of distributions $\mathcal{G}$ under a certain distribution discrepancy measure $\mathds{D}$. Then, the optimal parameters are selected for the resulting distribution. However, even with additional information about our contamination models, this “two-stage” procedure has two major drawbacks: the difficulty in choosing an appropriate family of distributions $\mathcal{G}$ and the inherent computational challenge to project probability distributions. Moreover, the first stage (projection) is usually sensitive to the choice of family $\mathcal{G}$. In contrast, we propose a novel approach that integrates outlier rectification (i.e., explicit projection) and parameter estimation into a joint optimization framework in (2.4). (ii) This min-min strategy has also been explored in Jiang & Xie (2024) using an artificially constructed rectification set. Their primary focus is to utilize the min-min formulation to recover well-studied robust statistics estimators. In contrast, our focus is on a new conceptual framework for formulating novel estimators which have not been studied before.
Illustrative Example.
Next, we aim to provide empirical evidence showcasing the potency and efficacy of the concave cost function on a toy example.
Figure 2: Visualization of contamination model: 55% are drawn with $x\sim\mathcal{N}(0,2)$ and $y=0.1-0.8x+0.2w$. 45% are drawn with $x\sim\mathcal{N}(4,2)$ and $y=10.1-0.8x+0.2w$. (a)-(c): In each case $w\sim N(0,1)$; The rectified data generated by the proposed statistically robust estimator with different budget $\delta$ and concave cost function (i.e., $r=0.5)$. (d)-(f): The rectified data generated by the proposed statistically robust estimator with different budget $\delta$ and convex cost function (i.e., $r=2)$.
In this example, we perform linear regression using our estimator with either a concave or a convex transport cost function. The estimation is performed on a sample of points in $\mathbb{R}^{2}$ drawn from a contaminated distribution (corresponding to $\mathbb{P}^{\prime}$) where 45% of the points are distributed according to an outlier distribution and 55% of the points are distributed according to the true uncontaminated distribution (corresponding to $\mathbb{P}_{\star}$). Our goal is to recover the line of best fit which corresponds to the uncontaminated distribution, which is $y=0.1-0.8x$. We illustrate the considerable differences which can occur when using each of these cost functions under various choices of $\delta$, the rectificatoin budget parameter. We depict this example in Figure 2.
The behavior of the esimtator using the concave cost function is depicted in Figures 2(a)-(c). Each of these subfigures shows the points drawn from each distribution, the rectified distribution, the true line of best fit for the uncontaminated distribution, and the line of best fit produced by our estimator, at various choices of $\delta$. In Figure 2(a), $\delta=0$, in Figure 2(b), $\delta=0.9$, and in Figure 2(a), $\delta=1.5$. As can be seen from the figures, as the budget $\delta$ increases, we note that the proposed estimator using the concave cost function becomes increasingly adept at mitigating the influence of outlier data. It achieves this by rectifying the outlying points, moving them into the bulk of the uncontaminated distribution. As a result, the line of best fit (depicted in red) aligns much more closely with the true line of best fit (depicted in black). It is able to perform this successful estimation because each outlier point can be moved far with a cheap cost (a “long haul”) due to the use of the concave transport cost function.
However, the convex cost function may result in suboptimal estimators, which can only move every data point (clean and contaminated) a little bit. Figures 2(d)-(f) display the behavior of the estimator using the convex cost function, which illustrates this effect. Each subfigure displays the rectified distribution as yellow-green points for a different setting of $\delta$ (i.e. (d) $\delta=0$, (e) $\delta=0.9$, and (f) $\delta=1.5$). As can be seen from the evolution of the rectified distribution, as $\delta$ increases, instead of the outlier points being moved towards the bulk of the distribution, all of the points move towards each other–even the points in the clean distribution. This is not ideal behavior, as the points from the clean distribution should stay in place. This defective behavior causes the line of best fit produced by the estimator using the convex cost function (which is depicted in red) fit to the rectified points to be severely affected by the outliers; instead of being close to the ground truth line of best fit (which is depicted in black), its slope is moved significantly upwards towards the outliers. This occurs because the convex cost function gives lower cost to smaller rectifications of the given data set, which is an inappropriate assumption for data sets containing outliers.
Appendices C.5 and C.6 depict the evolution of the rectified distribution as $\delta$ increases for the concave and convex cost functions, respectively. For our estimator which uses the concave cost, the evolution of the rectified distribution as $\delta$ increases shows that our estimator increasingly rectifies the points in order of their distance from the line of best fit in a predictable and stable way. In contrast, the convex estimator consistently moves all points (clean and corrupted) toward each other as $\delta$ changes, which achieves a poor estimate of the true line of best fit for all $\delta$ considered.

### v2

@@ Abstract
In this paper, we propose a novel conceptual framework to detect outliers using optimal transport with a concave cost function. Conventional outlier detection approaches typically use a two-stage procedure: first, outliers are detected and removed, and then estimation is performed on the cleaned data. However, this approach does not inform outlier removal with the estimation task, leaving room for improvement. To address this limitation, we propose an automatic outlier rectification mechanism that integrates rectification and estimation within a joint optimization framework. We take the first step to utilize the optimal transport distance with a concave cost function to construct a rectification set in the space of probability distributions. Then, we select the best distribution within the rectification set to perform the estimation task. Notably, the concave cost function we introduced in this paper is the key to making our estimator effectively identify the outlier during the optimization process. We demonstrate the effectiveness of our approach over conventional approaches in simulations and empirical analyses for mean estimation, least absolute regression, and the fitting of option implied volatility surfaces.
@@ 1 Introduction
The challenge of outlier removal has been documented for centuries: early work is introduced in e.g. (Gergonne, 1821) and (Peirce, 1852). Yet the outlier removal problem continues to interest practitioners and researchers alike due to the danger of distorted model estimation. A natural family of approaches followed in the literature takes the form of “two-stage” methods, which involve an outlier removal step followed by an estimation step. Methods within this family range from rules of thumb, such as removing outliers beyond a particular threshold based on a robust measure of scale like the interquartile range (Tukey, 1977), to various model-based or parametric distributional assumption-based tests (Thompson, 1985). While the two-stage approach can be useful by separating the estimation task from the outlier removal objective, it is not without potential pitfalls. For example, outlier detection which relies on a specific fitted model may tend to overfit to a particular type of outlier, which can mask the effect of other types (Rousseeuw and Leroy, 1987). Conversely, the outlier detection step may be overly conservative if it is not informed by the downstream estimation task; this can lead to a significant reduction in the number of observations available for model estimation in the next step, resulting in a loss of statistical efficiency (He and Portnoy, 1992).
@@
Robust statistics, pioneered by (Box, 1953), (Tukey, 1960), (Huber, 1964) and others such as (Hampel, 1968, 1971) offers alternative approaches for obtaining statistical estimators in the presence of outliers without removing them, particularly in parametric estimation problems such as linear regression. Beyond these, one closely related approach is the minimum distance functionals-based estimator, introduced by (Parr and Schucany, 1980; Millar, 1981; Donoho and Liu, 1988a, b; Park et al., 1995; Zhu et al., 2022; Jaenada et al., 2022). This method involves projecting the corrupted distribution onto a family of distributions using a distribution metric and selecting the optimal estimator for the resulting distribution. However, this projection mechanism is not informed by the estimation task (such as fitting a linear regression or neural network). Thus, without additional information about the contamination model or estimation task, it can be challenging to choose an appropriate family of distributions for projection, which may lead to limitations similar to those outlined above in practice.
@@
In this paper, we propose a novel approach to integrate both outlier detection and estimation in a joint optimization framework. A key observation is that statisticians aim to “clean” data before decisions are made; thus, ideal robust statistical estimators tend to be optimistic. To address this, we introduce the rectification set: a ball centered around the empirical (contaminated) distribution and defined by the optimal transport distance (see (Villani, 2009; Peyré et al., 2019)) in probability space. This rectification set aims to exclude potential outliers and capture the true underlying distribution, allowing us to minimize the expectation over the “best-case” scenarios, leading to a min-min problem. The study in Jiang and Xie (2024) also explores connections with min-min-type problems, but concentrates on the construction of artificially constructed rectification sets to cover certain existing estimators from the robust statistics literature. However, our primary focus is on introducing a novel rectification set, which is based on the optimal transport approach with a concave cost function.
@@
To automatically detect outliers during the estimation process, one of our main contributions is the use of a concave cost function for the optimal transport distance. This function encourages what we refer to as “long haul” transportation, in which the optimal transport plan moves only a portion of the data to a distant location, while leaving other parts unchanged. This strategic approach effectively repositions identified outliers closer to positions that better align with the clean data. Our novel formulation then involves minimizing the expected loss under the optimally rectified empirical distribution. This rectification is executed through the application of an optimal transport distance with a concave cost function, thereby correcting outliers to enhance performance within a fixed estimation task. More importantly, our method distinguishes itself from distributionally robust optimization (DRO) (Ben-Tal et al., 2013; Bayraksan and Love, 2015; Wiesemann et al., 2014; Delage and Ye, 2010; Gao and Kleywegt, 2022; Blanchet and Kang, 2017; Shafieezadeh Abadeh et al., 2015; Shafieezadeh-Abadeh et al., 2019; Sinha et al., 2018; Kuhn et al., 2019) due to the distinctive correction mechanisms initiated by the robust formulation employing a min-min strategy.
@@
As we discuss in Section 2 below, the timing of error generating differs crucially between DRO and robust statistics. DRO approach employs a min-max game strategy to control the worst-case loss over potential post-decision distributional shifts. In contrast, the robust estimator acts after the pre-decision distributional contamination materializes. Thus the approach of robust statistics can be motivated as being closer to a max-min game against nature. As a consequence, in robust statistics, the adversary moves first, and therefore the statistician can be more optimistic that they can rectify the contamination applied by the nature thus motivating the min-min strategy suggested above. However, it is also worth mentioning that (Kang and Bansal, 2023; Nietert et al., 2024) are also trying to formulate the outlier-robust problem in a min-max form, incorporating additional information about the contamination model to further shrink the ambiguity set.
@@
(i) Novel statistically robust estimator. We propose a new statistically robust estimator that incorporates a novel rectification set constructed using the optimal transport distance. By employing a concave cost function within the optimal transport distance, our estimator enables automatic outlier rectification. We prove that the optimal rectified distribution can be found via a finite convex program, and show that we can determine the optimal rectification set uncertainty budget via cross-validation.
(ii) Connection to adaptive quantile regression. For mean estimation and least absolute regression, we demonstrate that our robust estimator is equivalent to an adaptive quantile (regression) estimator, with the quantile controlled by the budget parameter $\delta$. Furthermore, we prove that the optimal rectified distribution exhibits a long haul structure that facilitates outlier detection.
(iii) Effectiveness in estimating the option implied volatility surface. We evaluate our estimator on various tasks, including mean estimation, least absolute regression, and option implied volatility surface estimation. Our experimental results demonstrate that our estimator produces surfaces that are 30.4% smoother compared to baseline estimators, indicating success in outlier removal, as surfaces should be smooth based on structural financial properties. Additionally, it achieves an average reduction of 6.3% in mean average percent error (MAPE) across all estimated surfaces, providing empirical evidence of the effectiveness of our rectifying optimal transporter.
@@ 2 DRO and Robust Statistics as Adversarial Games
In this section, we summarize conceptually how robust statistics is different from DRO (for more details, we refer the reader to e.g. Blanchet et al. (2024)). To lay a solid mathematical foundation, we begin by investigating a generic stochastic optimization problem. Here, we assume that $Z$ is a random vector in a space $\mathcal{Z}\subseteq\mathbb{R}^{d}$ that follows the distribution $\mathbb{P}_{\star}$. The set of feasible model parameters is denoted $\Theta$ (assumed to be finite-dimensional to simplify). Given a realization $z$ and a model parameter $\theta\in\Theta$ the corresponding loss is $\ell(\theta,z)$. A standard expected loss minimization decision rule is obtained by solving
$\min_{\theta\in\Theta}~{}\mathbb{E}_{\mathbb{P}_{\star}}[\ell(\theta,\xi)]=\int_{\Xi}\ell(\theta,\xi)~{}\mathrm{d}\mathbb{P}_{\star}(\xi).\vspace{-5pt}$ (1)
Since $\mathbb{P}_{\star}$ is generally unknown, to approximate the objective function in (1), we often gather $n$ independent and identically distributed (i.i.d.) samples $\{z_{i}\}_{i=1}^{n}$ from the unknown data-generating distribution $\mathbb{P}_{\star}$ and consider the empirical risk minimization counterpart,
$\min_{\theta\in\Theta}~{}\mathbb{E}_{\hat{\mathbb{P}}_{n}}[\ell(\theta,Z)]=\frac{1}{n}\sum_{i=1}^{n}\ell(\theta,z_{i}),\vspace{-5pt}$ (2)
where ^$\hat{\mathbb{P}}_{n}$ denotes the empirical measure $\frac{1}{n}\sum_{i=1}^{n}\delta_{z_{i}}$ and $\delta_{z}$ is the Dirac measure centered at $z$. These problems are solved within the context of the general data-driven decision-making cycle below:
Contaminated data-generating distribution Data-driven model Out-of-sample environment $\mathbb{P}_{\star}$$\ {\mathbb{P}^{\prime}}$$\ \ {\mathbb{P}}_{n}^{\prime}$$\ \ \widetilde{\mathbb{P}}$
@@
In the cycle depicted in Figure 1, we usually collect $n$ i.i.d samples from the unknown data-generating distribution $\mathbb{P}^{\prime}$, which may be identical to or distinct from the clean distribution $\mathbb{P}_{\star}$. Subsequently, we make a decision (e.g., parameter estimation) based on a model, ${\mathbb{P}}_{n}^{\prime}$, built from these samples. Such a model could be parametric or non-parametric. These decisions are then put into action within the out-of-sample environment $\tilde{\mathbb{P}}$, which may or may not conform to the distribution $\mathbb{P}_{\star}$. In this general cycle, the sample average method (2) may lead to poor out-of-sample guarantees. This motivates the use of alternative approaches. In the following paragraphs, we describe two of these approaches, DRO and robust statistics, by treating them as adversarial games. The crucial distinction lies in the timing of the contamination or attacks.
(i) (DRO: $\widetilde{\mathbb{P}}\neq{\mathbb{P}}_{\star}={\mathbb{P}^{\prime}}$) The attack occurs in the post-decision stage. In this scenario, the out-of-sample environment $\tilde{\mathbb{P}}$ diverges from the data-generating distribution $\mathbb{P}_{\star}$, and no contamination occurs before the decision, implying that $\mathbb{P}_{\star}=\mathbb{P}^{\prime}$. For example, in adversarial deployment scenarios, malicious actors can deliberately manipulate the data distribution to compromise the performance of trained models. With full access to our trained machine learning model, the adversary endeavors to create adversarial examples specifically designed to provoke errors in the model’s predictions.
To ensure good performance in terms of the optimal expected population loss over the out-of-sample distribution, the DRO framework introduces an uncertainty set $\mathcal{B}({\mathbb{P}}^{\prime}_{n})$ to encompass discrepancies between the in-sample-distribution ${\mathbb{P}}_{n}^{\prime}$ and the out-of-sample distribution $\tilde{\mathbb{P}}$. Subsequently, the DRO formulation minimizes the worst-case loss within this uncertainty set, thereby aiming to solve
$\min_{\theta\in\Theta}\sup_{\mathbb{Q}\in\mathcal{B}(\mathbb{P}_{n}^{\prime})}\mathbb{E}_{\mathbb{Q}}[\ell(\theta,Z)].\vspace{-5pt}$ (3)
@@
$\min_{\theta\in\Theta}\min_{\mathbb{Q}\in\mathcal{R}({\mathbb{P}}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}\left[\ell(\theta,Z)\right].$ (4)
Our goal in this paper is to develop (3), which is completely distinct from DRO.
@@ 3 Automatic Outlier Rectification Mechanism
We now introduce our primary contribution by delineating (3). Our estimator is crafted to incorporate outlier rectification and parameter estimation within a unified optimization framework, facilitating automatic outlier rectification. A natural question arises from (3): how do we construct the rectification set in the space of probability distributions to correct the observed data set? In this paper, we employ an optimal transport distance to create a ball centered at the contaminated empirical distribution $\mathbb{P}_{n}^{\prime}$.
Definition 3.1 (Rectification Set).
@@
$\mathcal{R}(\mathbb{P}_{n}^{\prime})=\{\mathbb{Q}\in\mathcal{P}(\mathcal{Z}):\mathds{D}_{c}(\mathbb{Q},\mathbb{P}^{\prime}_{n})\leqslant\delta\},$ (5)
where $\delta>0$ is a radius, and $\mathds{D}_{c}$ is a specific optimal transport distance defined in Definition 3.2.
Definition 3.2.
(Optimal Transport Distance (Peyré et al., 2019; Villani, 2009)) Suppose that $c(\cdot,\cdot):\mathcal{Z}\times\mathcal{Z}\rightarrow[0,\infty]$ is a lower semi-continuous cost function such that $c(z,z^{\prime})=0$ for all $z,z^{\prime}\in\mathcal{Z}$ satisfying $z=z^{\prime}$. The optimal transport distance between two probability measures $\mathbb{P},\mathbb{Q}\in\mathcal{P}(\mathcal{Z})$ is
@@
If we consider $c(z,z^{\prime})=\|z-z^{\prime}\|^{r}$ as a metric defined on $\mathbb{R}^{d}$, where $r\in[1,\infty)$, then the distance metric $\mathds{D}_{c}^{1/r}(\mathbb{P},\mathbb{Q})$ corresponds to the $r$-th order Wasserstein distance (villani2009optimalWasserstein). In this paper, we pioneer the utilization of concave cost functions, exemplified by $c(z,z^{\prime})=\|z-z^{\prime}\|^{r}$ where $r\in(0,1)$ in statistical robust estimation. We note that $\mathds{D}_{c}(\mathbb{P},\mathbb{Q})$ (as opposed to $\mathds{D}_{c}^{1/r}(\mathbb{P},\mathbb{Q})$ ) is also a metric for $r\in[0,1)$. The rationale behind selecting a concave cost function is intuitive: it promotes what we colloquially refer to as long haul transportation plans, enabling outliers to be automatically moved significant distances back towards the central tendency of data distribution. This, in turn, facilitates automatic outlier rectification. Concave costs promote long hauls due to their characteristic of exhibiting decreasing marginal increases in transportation cost. In other words, if an adversary decides to transport a data point by $\|\Delta\|$ units, it becomes cheaper to continue transporting the same point by an additional $\varepsilon$ distance compared to moving another point from its initial location. We make further remarks about the connection between our estimator and prior work in Appendix B.1.
Illustrative Example. We provide empirical evidence showcasing the efficacy of the concave cost function on simulated data in Figure 2 and 3. This example shows that our concave cost function is critical for moving the outliers (orange) properly to the bulk of the clean distribution (blue). For the concave cost, only the outliers are rectified (green), resulting in the proper line of best fit.
(a) $\delta=0$
(b) $\delta=0.9$
(c) $\delta=1.5$
Figure 2: The rectified data generated by our estimator with concave cost function ($r=0.5)$.
(a) $\delta=0$
(b) $\delta=0.9$
(c) $\delta=1.5$
Figure 3: The rectified data generated by our estimator with convex cost function ($r=2)$.
However, for the convex cost, regardless of the budget, all points both clean and corrupted are rectified towards each other instead, which results in an incorrect line of best fit. This illustrates succinctly the importance of our novel concave cost. More details on this example are given in Appendix F.1.